# Seer Self-Consistency: Advance Budget Estimation for Adaptive Test-Time Scaling

### v1

@@ 1 Introduction
Large Language Models (LLMs) (Hurst et al., 2024; Yang et al., 2024; Guo et al., 2025) have demonstrated exceptional performance in reasoning tasks, particularly in mathematics. By leveraging methods such as self-consistency (Wang et al., 2022), LLMs can substantially improve their performance during inference through increased computation (Chen et al., 2025; Zhang et al., 2025) without additional training. While improving performance, Test-Time Scaling (TTS) methods also significantly increase the computational cost (Brown et al., 2024; Chen et al., 2024a). Many efforts (Feng et al., 2025; Liu et al., 2025) focus on optimizing the trade-off between performance and efficiency.
@@
For ensemble methods like self-consistency, a critical challenge is dynamically determining the optimal number of ensemble samples based on the complexity of the input instance (Wang et al., 2024). AC (Aggarwal et al., 2023) generates responses sequentially and decides whether to terminate scaling based on the diversity of the current answers and a specified threshold. ESC (Li et al., 2024) employs a sliding window to track the diversity of the current responses, enabling dynamic early stopping for self-consistency. By dynamically adjusting the sampling budget, these methods significantly reduce token consumption during the TTS phase.
@@ 3.2 System 1 for Estimation
As illustrated in Figure 3, given a problem $X$, System 1 first performs direct answer generation with a sampling size of $M$, producing an answer set $\mathrm{Ans}=\{A_{1},A_{2},\dots,A_{M}\}$.
@@ 3.3 System 2 for Scaling
This empirical setting provides a practical balance between efficiency and accuracy.
@@ 4.1 Experimental Setup
We conduct evaluations on the following datasets. MATH500 (Lightman et al., 2023) represents a representative subset of problems from the MATH dataset (Hendrycks et al., 2021). AIME2024 (AoPS, 2024), AIME2025 (AoPS, 2025), and AMC2023 (AoPS, 2023) consist of high-difficulty mathematical competition problems. GPQA Diamond (Rein et al., 2024) serves as a benchmark for graduate-level STEM reasoning tasks, providing a challenging evaluation setting. We evaluate on the following models: DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025), and Qwen3-4B (Yang et al., 2025). The first two are representative reasoning models, while Qwen3-4B supports both thinking and non-thinking modes, and we conduct evaluations under both settings.
@@ 4.2 Main Results
Across all models and datasets, SeerSC consistently achieves accuracy comparable to the baseline methods. On medium-difficulty tasks such as MATH500 and AMC2023, its accuracy is essentially on par with SC, with only minor fluctuations (within ±1 percentage point). On the more challenging datasets AIME2024/2025 and GPQA-Diamond, BoX also sustains stable performance without noticeable degradation. Given the stochasticity introduced by high-temperature decoding, the overall accuracy level remains comparable to that of the baselines.
@@ 5.1 Latency Analysis
As shown in Figure 4, the latency of System 1 accounts for less than 1% of the total latency of SeerSC. In practical applications, the overhead of System 1 would be even smaller, since the KV cache produced during its prefilling stage can be reused in System 2. As a result, the additional delay corresponds only to decoding a few tokens and is therefore negligible.
@@ 5.3 Orthogonality
It is worth noting that SeerSC is orthogonal to several existing parallel test-time scaling optimization techniques, such as weighted voting and path pruning. We conducted experiments to validate this claim.
@@
Existing weighted vote methods typically assign weights to reasoning paths based on their confidence scores, such that high-confidence paths exert greater influence in the majority voting process, thereby improving voting accuracy (Wan et al., 2025; Zhou et al., 2025). Building on standard SeerSC, we incorporated the weighted voting strategy from (Fu et al., 2025), which leverages tail confidence—the confidence scores of the last window in a path—to determine its weight in voting. As shown in Table 3, both DeepSeek-R1-Distill-Qwen-7B and Qwen3-4B (non-thinking) achieved accuracy improvements across datasets. Notably, on AIME-2024, Qwen3-4B (non-thinking) outperformed vanilla SeerSC by +3 percentage points. Moreover, combining the two methods does not affect token consumption or latency efficiency.
@@ A.1 Optimal $N$ for SC
Table 4: Optimal $N$ for SC on different models and datasets.
Model MATH500 AIME2024 AIME2025 AMC2023 GPQA-D
DeepSeek-R1-Distill-Qwen-7B 8 32 32 8 16
DeepSeek-R1-Distill-Llama-8B 8 32 32 8 8
Qwen3-4B (thinking) 8 16 16 8 8
Qwen3-4B (non-thinking) 8 32 16 8 16

### v2

@@ 1 Introduction
Large Language Models (LLMs) (Hurst et al., 2024; Yang et al., 2024; Guo et al., 2025) have demonstrated exceptional performance in reasoning tasks, particularly in mathematics. By leveraging methods such as self-consistency (Wang et al., 2022), LLMs can substantially improve their performance during inference through increased computation (Chen et al., 2025; Zhang et al., 2025) without additional training. While improving performance, Test-Time Scaling (TTS) methods also significantly increase the computational cost (Brown et al., 2024; Chen et al., 2024a). Many efforts (Feng et al., 2025; Liu et al., 2025b) focus on optimizing the trade-off between performance and efficiency.
@@
For ensemble methods like Self-Consistency (SC), a critical challenge is dynamically determining the optimal number of ensemble samples based on the complexity of the input instance (Wang et al., 2024). Adaptive Consistency (AC) (Aggarwal et al., 2023) generates responses sequentially and decides whether to terminate scaling based on the diversity of the current answers and a specified threshold. Early-Stopping Self-Consistency (ESC) (Li et al., 2024) employs a sliding window to track the diversity of the current responses, enabling dynamic early stopping for self-consistency. By dynamically adjusting the sampling budget, these methods significantly reduce token consumption during the TTS phase.
@@ 3.2 System 1 for Estimation
As illustrated in Figure 3, given a problem $X$, System 1 first performs direct answer generation (see Appendix A.3 for implementation details on reasoning models) with a sampling size of $M$, producing an answer set $\mathrm{Ans}=\{A_{1},A_{2},\dots,A_{M}\}$.
@@ 3.3 System 2 for Scaling
This empirical setting provides a practical balance between efficiency and accuracy (see Appendix B for a detailed sensitivity analysis).
@@ 4.1 Experimental Setup
We conduct evaluations on the following datasets. MATH500 (Lightman et al., 2023) represents a representative subset of problems from the MATH dataset (Hendrycks et al., 2021). AIME2024 (AoPS, 2024), AIME2025 (AoPS, 2025), and AMC2023 (AoPS, 2023) consist of high-difficulty mathematical competition problems. GPQA Diamond (Rein et al., 2024) serves as a benchmark for graduate-level STEM reasoning tasks, providing a challenging evaluation setting. We evaluate on the following models: DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025), and Qwen3-4B (Yang et al., 2025). The first two are representative reasoning models, while Qwen3-4B supports both thinking and non-thinking modes, and we conduct evaluations under both settings. Additionally, we extend our evaluation to the larger-scale QwQ-32B (Team, 2025) model to verify generalizability (see Appendix C.1).
@@ 4.2 Main Results
Across all models and datasets, SeerSC consistently achieves accuracy comparable to the baseline methods. On medium-difficulty tasks such as MATH500 and AMC2023, its accuracy is essentially on par with SC, with only minor fluctuations (within ±1 percentage point). On the more challenging datasets AIME2024/2025 and GPQA-Diamond, SeerSC also sustains stable performance without noticeable degradation. Given the stochasticity introduced by high-temperature decoding, the overall accuracy level remains comparable to that of the baselines.
@@ 5.1 Latency Analysis
As shown in Figure 4, the latency of System 1 accounts for less than 1% of the total latency of SeerSC. In practical applications, the overhead of System 1 would be even smaller, since the KV cache produced during its prefilling stage can be reused in System 2. As a result, the additional delay corresponds only to decoding a few tokens and is therefore negligible (see Appendix C.2 for a evaluation of Time to First Token).
@@ 5.3 Orthogonality
It is worth noting that SeerSC is orthogonal to several existing parallel test-time scaling optimization techniques, such as weighted voting and path pruning. We conducted experiments to validate this claim. For extended experiments regarding orthogonality, please refer to Appendix C.3 and C.4.
@@
Existing weighted voting assigns confidence-based weights to paths, allowing high-confidence samples to exert greater influence and improve accuracy. (Wan et al., 2025; Zhou et al., 2025). Building on standard SeerSC, we incorporated the weighted voting strategy from (Fu et al., 2025), which leverages tail confidence—the confidence scores of the last window in a path—to determine its weight in voting. As shown in Table 3, both DeepSeek-R1-Distill-Qwen-7B and Qwen3-4B (non-thinking) achieved accuracy improvements across datasets. Notably, on AIME-2024, Qwen3-4B (non-thinking) outperformed vanilla SeerSC by +3 percentage points. Moreover, combining the two methods does not affect token consumption or latency efficiency.
@@ A.2 Implementation details of SeerSC
A.3 System 1 Implementation for Reasoning Models
Reasoning models (e.g., DeepSeek-R1 series or QwQ) typically generate an extended reasoning process enclosed within <think> and </think> tags before producing the final answer. However, System 1 requires rapid, direct answer generation to estimate entropy efficiently.
To enable this, we employ a forced decoding strategy. Specifically, we append a sequence to the model input that simulates the completion of the thinking process and prompts the immediate start of the final answer. The appended sequence is:
[…]
Example.
Given the input query:
[…]
A standard generation would produce:
[…]
With our suffix injection, the effective input becomes:
[…]
And the model output becomes:
[…]
This technique ensures that System 1 inference remains computationally lightweight and fast.
Table 6: Experimental results on QwQ-32B model. SeerSC demonstrates strong generalizability, achieving significant reductions in token usage and latency on diverse datasets, while maintaining accuracy comparable to baselines. We report accuracy (%), average tokens ($\times 10^{3}$), and average latency (s).
Methods MATH500 AIME2024 AIME2025 AMC2023 GPQA-D
[…]
QwQ-32B
CoT 91.1 65.1 51.8 81.1 63.5
SC 93.6 33.3 16.2 74.0 93.4 95.3 66.0 101.3 87.6 86.0 51.7 29.6 66.4 69.6 58.0
AC 93.5 21.0 19.1 73.3 73.6 135.2 65.3 91.4 123.7 86.0 36.9 41.2 66.1 53.9 70.5
ESC 93.4 23.9 22.2 73.6 78.5 127.1 65.1 93.8 115.5 86.5 40.4 45.8 66.1 58.6 77.9
SeerSC 93.0 19.5 9.6 74.4 88.3 90.6 66.2 95.4 83.2 85.5 36.7 19.1 66.3 59.7 43.4
Appendix B Hyperparameter Sensitivity Analysis
The entropy thresholds $\tau_{1}$ and $\tau_{2}$ in Equation 5 are critical hyperparameters that determine the trade-off between accuracy and latency. We empirically determined these values based on performance on the MATH training dataset using the DeepSeek-R1-Distill-Qwen-7B model.
Impact of the First Threshold ($\tau_{1}$).
We first investigated the impact of a single threshold $\tau_{1}$ (effectively setting $\tau_{2}=\tau_{1}$, resulting in a binary budget allocation of either 1 or $N$). As shown in Table 5(a), increasing $\tau_{1}$ leads to more samples being assigned a budget of 1. Consequently, the method transitions from the behavior of SC (high accuracy, high latency) to that of CoT (low accuracy, low latency). A threshold of $1/10$ provides a strong initial filtering, significantly reducing latency while preserving most of the accuracy gains.
Impact of the Second Threshold ($\tau_{2}$).
To achieve a finer-grained budget allocation, we introduced the second threshold $\tau_{2}$ (assigning a budget of $N/2$ for samples with entropy between $\tau_{1}$ and $\tau_{2}$). We fixed $\tau_{1}=1/10$ and varied $\tau_{2}$. As presented in Table 5(b), introducing $\tau_{2}$ further optimizes the efficiency. The setting $(\tau_{1}=1/10,\tau_{2}=1/3)$ achieves a significant reduction in latency (from 7.4s to 5.3s) compared to using a single threshold, with only a negligible drop in accuracy (0.3%).
Summary.
These empirical results justify our choice of $\tau_{1}=\frac{1}{10}\log M$ and $\tau_{2}=\frac{1}{3}\log M$. While this configuration may not be the global optimum for every specific scenario, our extensive experiments across diverse datasets and model scales (ranging from 4B to 32B) demonstrate its consistent effectiveness and robustness. We adopt these fixed values to avoid the additional computational overhead associated with per-task hyperparameter searching on validation sets. The exploration of more fine-grained or adaptive thresholding mechanisms remains a promising avenue for our future work.
Appendix C Additional Experiments
C.1 Generalizability to Larger Models
To assess generalizability, we extended our evaluation to the larger-scale QwQ-32B model with a sampling budget of $N=8$. As shown in Table 6, the core conclusions observed in smaller models remain valid: SeerSC maintains accuracy parity with the SC baseline while significantly reducing computational costs. This confirms that SeerSC’s adaptive budget allocation remains robust and effective at larger model scales.
C.2 Impact on Time to First Token (TTFT)
We further investigated the impact of the System 1 pre-estimation stage on the Time to First Token (TTFT), a critical metric for streaming applications. We measured the TTFT on a single AIME2024 query using DeepSeek-R1-Distill-Qwen-7B served by vLLM.
As shown in Table 7, although SeerSC requires a pre-estimation step (generating 64 direct answers), the increase in TTFT is minimal. Specifically, the TTFT increases from 0.022s (SC) to 0.234s (SeerSC), adding an overhead of only $\approx 0.2$s. Given that the average completion time for a complex reasoning task (e.g., AIME) is typically around 30 seconds, this slight delay is negligible relative to the total inference latency and does not perceptibly affect the user experience.
Table 7: Comparison of Time to First Token (TTFT) between SC and SeerSC on a single AIME query. The overhead introduced by System 1 is negligible ($\approx 0.2$s).
Method TTFT (s) Overhead (s)
SC 0.022
SeerSC 0.234 +0.212
Table 8: Impact of Weighted Voting on baselines (DeepSeek-R1-Distill-Qwen-7B). Values in parentheses denote changes relative to vanilla versions. Despite accuracy gains in baselines, SeerSC retains superior latency efficiency.
Methods MATH500 AMC2023 GPQA-D
[…]
SC 93.9 8.4 91.6 13.2 48.1 51.8
+ Weighted Vote 94.1 (+0.2) 8.4 92.5 (+0.9) 13.2 48.7 (+0.6) 52.1
AC 93.8 12.4 91.3 24.8 47.7 73.1
+ Weighted Vote 94.0 (+0.2) 12.5 92.3 (+1.0) 25.2 47.7 (+0.0) 73.5
ESC 93.6 13.9 91.1 26.7 48.0 82.0
+ Weighted Vote 93.4 (-0.2) 13.8 92.1 (+1.0) 26.7 48.6 (+0.6) 82.3
SeerSC 93.0 5.3 91.5 9.4 48.6 37.8
+ Weighted Vote 93.8 (+0.8) 5.3 92.2 (+0.7) 9.4 49.0 (+0.4) 38.1
Table 9: Experimental results of combining SeerSC with other TTS paradigms. The results demonstrate that SeerSC is compatible with both reward model-based and length-controlled methods.
(a) Combination with Best-of-N (Generator: DeepSeek-R1-Distill-Qwen-7B; RM: Skywork-Reward-V2-Qwen3-8B). BoN improves accuracy, at the cost of scoring latency.
Method MATH500 GPQA-D
Acc. (%) Token ($10^{3}$) Lat. (s) Acc. (%) Token ($10^{3}$) Lat. (s)
SeerSC 93.0 17.7 5.3 48.6 95.3 37.8
+ Best-of-N 93.4 (+0.4) 17.7 7.3 51.0 (+2.4) 95.3 48.7
(b) Combination with TokenSkip (Model: Qwen2.5-7B-Instruct fine-tuned with TokenSkip). TokenSkip reduces token usage significantly, demonstrating orthogonality with SeerSC.
Method GSM8K MATH500
Acc. (%) Token Lat. (s) Acc. (%) Token Lat. (s)
SeerSC 92.9 1437 0.8 78.6 2726 0.5
+ TokenSkip 91.0 (-1.9) 1205 (-16%) 0.8 76.0 (-2.6) 2379 (-13%) 0.5
C.3 Impact of Weighted Voting on Baselines
We extend the confidence-based weighted voting (Fu et al., 2025) to baselines (SC, AC, ESC) on DeepSeek-R1-Distill-Qwen-7B to evaluate potential gains. As shown in Table 8, while weighted voting yields marginal accuracy improvements for baselines, it fails to address their inherent latency bottlenecks. In contrast, SeerSC maintains a significant efficiency advantage—e.g., on GPQA-D, it is $1.4\times$ faster than SC and $\approx 2\times$ faster than AC/ESC—confirming that our efficiency gains stem from proactive budget allocation rather than the voting mechanism.
C.4 Compatibility with Other TTS Paradigms
In Section 5.3, we demonstrated that SeerSC is orthogonal to parallel optimization techniques such as weighted voting and path pruning. Beyond these strategies, SeerSC serves as a flexible budget allocation framework that can also be integrated with broader Test-Time Scaling (TTS) paradigms. To demonstrate this extensibility, we conducted experiments combining SeerSC with Reward Model-based selection (Best-of-N) and Length-Controlled scaling (TokenSkip).
Combining with Best-of-N.
We integrated SeerSC with Best-of-N (BoN) using DeepSeek-R1-Distill-Qwen-7B as the generator and Skywork-Reward-V2-Qwen3-8B (Liu et al., 2025a) as the reward model. Instead of majority voting, we used the reward model to rank the candidates generated within the SeerSC budget.
As shown in Table 9(a), this combination yields consistent accuracy improvements, particularly on the challenging GPQA-D benchmark (+2.4%). However, the introduction of a reward model incurs additional latency for scoring, which is a trade-off users must consider.
Combining with Length-Controlled Scaling.
We also explored combining SeerSC with TokenSkip (Xia et al., 2025), a sequential scaling method designed to reduce token usage for reasoning models. We used a Qwen2.5-7B-Instruct model fine-tuned with TokenSkip (length parameter 0.7).
As shown in Table 9(b), combining SeerSC with TokenSkip significantly reduces token consumption (up to 16% on GSM8K). This comes with a slight accuracy drop, which is primarily attributable to the performance degradation inherent in the TokenSkip fine-tuning process.